OpenStudy (anonymous):

does the system have one solution, no solution, or an infinite number of solutions? 5x-y=10 x+6y=33 a. one b. infinite numbers c. none

OpenStudy (anonymous):

Muitiply(6) : 30x-6y=60 x+6y=33 Add: 31x+oy=93 x =93/31 =3 subtitution for y: 3+6y=33 y=(33-3)/6 y=5 yes, solution